OpenStudy (magbak):

PLEASE HELP ME!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! I WILL AWARD MEDAL!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Simplify the following equation and find the restriction y^2 -y -6 -------- y^2 -4

OpenStudy (magbak):

@Mertsj @zscdragon @Mousam

OpenStudy (anonymous):

same thing factorise the top and the bottome and cancel out the similar terms

OpenStudy (magbak):

Ok it is y-3 ---- y-2 But the restriction is the problem.

OpenStudy (agent0smith):

You mean the restricted values? They'll be values that make the denominator equal to zero (you can't divide by zero) You need to solve for y \[\Large y^2 - 4 = 0\]

OpenStudy (magbak):

I know that it is just I got it wrong last tim.

OpenStudy (agent0smith):

Can you factor y^2-4? you should get \[\large (y+2)(y-2) = 0\]Now solve that.

OpenStudy (magbak):

Oh so the restriction is pluse or minus 2

OpenStudy (anonymous):

and z =1 too

OpenStudy (magbak):

Thank you so much so I am right,

OpenStudy (agent0smith):

Er,, there's no z?

OpenStudy (magbak):

yes no Z

OpenStudy (anonymous):

where z = y^2 -y -6 -------- y^2 -4

OpenStudy (agent0smith):

Restricted values are concerned with the domain, not range.

OpenStudy (anonymous):

trust me or if you font trust me.. try a graphic calculator

OpenStudy (agent0smith):

And there's no z. The equation at the start was y^2 -y -6 -------- y^2 -4

OpenStudy (anonymous):

ok...

OpenStudy (agent0smith):

z=1 is not a restricted value even if z = y^2 -y -6 -------- y^2 -4 Restricted values are values in the domain, not range.

OpenStudy (magbak):

I still have a few more questions any one willing to help me.

OpenStudy (agent0smith):

@Mousam you're talking about a horizontal asymptote. Not a restricted value.

OpenStudy (anonymous):

i thought \[z=\frac{ y-3 }{ y-2 }\]\[\lim_{y \rightarrow \infty} z=\frac{ 1-\frac{ 3 }{ y } }{ 1-\frac{ 2 }{ y } } = \]

OpenStudy (anonymous):

oh so he is just asking for the vertical ones?

OpenStudy (agent0smith):

Again, restricted values are values not in the domain. y=2 and y=-2 are not in the domain. y=2 is also a vertical asymptote. y=-2 is a point of discontinuity. z=1 is a horizontal asymptote.